Write an equation of the line that passes through the point (18, −7) and is perpendicular to the line y = −9x + 8.
Answer by mananth(16946)

Write an equation of the line that passes through the point (18, -7) and is perpendicular to the line y = -9x + 8.
...
(18, -7) and is perpendicular to the line y = -9x + 8.
...
y = -9x + 8.
...
compare with y= mx +b
..
m= -9 the slope
The perpendicular line will have slope 1/9.
....
(18, -7) and m= 1/9
Find b by plugging (18, -7) m= 1/9 ,
y = mx + b
-7=18*1/9 +b
-7=2+b
b=-9
...
The equation will be
y = (x/9)-9
...
m.ananth@hotmail.ca